# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 7 Conclusion
In this work, we have investigated the cascades of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence. The study is based on direct numerical simulations of a dilute polymeric solution at $Re_{\lambda}\approx 460$, with $1/9\leq De\leq 9$. We have extended the formulation introduced by Baj et al. (2022) and derived the exact scale-by-scale budget equations for the energy, enstrophy and helicity, $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, for polymeric turbulent flows. The equations are general and remain valid for a generic inhomogeneous and anisotropic turbulent flow as well. They capture the mechanisms of production, transfer and dissipation in the combined space of scales and positions. Compared to purely Newtonian flows, the polymer/fluid interaction introduces additional sink/source process and an alternative scale-space transfer for the three quantities. The newly derived equations have been then tailored to homogeneous and isotropic turbulence for the present problem.
@@
For $De\geq 1$ polymers effectively modify the cascades of energy and helicity, the two inviscid invariants of the $3D$ Navier–Stokes equations. Both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ are injected in the system at the largest scales and transferred to the small scales by two distinct transfer processes driven by fluid inertia and fluid/polymer interactions, where they are dissipated away. At large scales, the nonlinear cascade dominates, while at small scales the polymer-driven transfer takes over. The cross-over scale between the two dominant processes agrees fairly well $r_{p}^{*}$ at which $\tau_{p}/\tau_{f}(r_{p}^{*})=1$, where $\tau_{p}$ and $\tau_{f}(r)$ are the polymer relaxation time and the characteristic turnover time of eddies at scale $r$. When $r<r_{p}^{*}$ and $\tau_{p}>\tau_{f}(r)$ polymers effectively interact with the $r$-fluctuations of the carrier flow. The coexistence of the two transfer mechanisms leads to a multiscaling behaviour of the velocity and helicity structure functions. At small $De$ and large $r$, $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ exhibit the same power law predicted by the Kolmogorov theory, i.e. $\left\langle{\delta q^{2}}\right\rangle\sim\left\langle{\delta h}\right\rangle\sim r^{2/3}$ with $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle\approx 0.75$, consistently with helicity behaving as a passive scalar in HIT (Borue & Orszag, 1997). For $De\geq 1$ and at smaller scales, instead, both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ deviate from the Kolmogorov predictions, with the latter showing a steeper slope. For $De=1$ we measure $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ (in agreement with Zhang et al., 2021) and $\left\langle{\delta h}\right\rangle\sim r^{0.9}$. Accordingly, we observe that compared to the purely Newtonian case, the relative helicity $r\left\langle{\delta h}\right\rangle/2\left\langle{\delta q^{2}}\right\rangle$ increases with $De$ at all scales, indicating that polymers favour events that break mirror symmetry at small scales.
@@
A closer look at the energy fluxes reveals that polymers deplete the nonlinear cascade by weakening both the direct and inverse extreme events of both cascades, leading to a less skewed flux distribution as $De$ increases. For $r\gtrapprox r_{p}^{*}$, the nonlinear flux distribution collapses reasonably well for all $De$ once the quantities are normalised with their standard deviation. Moving to the polymeric flux, we observe that the amount of energy carried by the polymer-driven cascade increases with $De$ at all scales. Similar to the nonlinear flux, on average the polymeric flux transfers energy from larger to smaller scales, but local events with intense backscatter exist which transfer energy from small to large scales. However, the probability of inverse transfer decreases as $De$ increases.
@@
Unlike energy and helicity, enstrophy is not an inviscid invariant of the $3D$ Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, which is related to the enstrophy produced by vortex stretching-like processes (Davidson, 2004). The conventional notion of cascade used for energy and helicity, thus, does not apply in this case, as most of the enstrophy is directly generated at the small scales. However, vortex stretching is active in the inertial range which results in an increasing rate of transfer of $\left\langle{\delta\omega^{2}}\right\rangle$ at small $r$, being maximum at the end of the inertial range. For $De\geq 1$, the transfer via the polymeric route is comparable with the fluid nonlinear contribution and, unlike for $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, there is no transition with the two mechanisms coexisting in the entire inertial range. Overall, the net effect of polymers is to largely weaken vortex stretching. At large $De$, particularly, viscous dissipation balances the enstrophy generation due to fluid/polymer interaction. The underlying local flow topology further reveals that vortex stretching modulation is a result of polymers promoting events with a two-dimensional state of straining, like shear and planar extensional flows. Accordingly, we observe that polymers favour events where rotation and stretching are equally strong (being typical of vortex sheets and shear layers), rather than those dominated by rigid rotation.
@@ Acknowledgments
The authors acknowledge the computer time provided by the Scientific Computing and Data Analysis section of Research Support Division at OIST, and by HPCI, under the Research Project grants hp210269, hp220099, and hp230018.

### v2

@@ 6 Conclusion
In this work, we have investigated the cascades of energy, enstrophy and helicity in homogeneous and isotropic polymeric turbulence. The study is based on direct numerical simulations of a dilute polymeric solution at $Re_{\lambda}\approx 460$, with $1/9\leq De\leq 9$. We have extended the formulation introduced by Baj et al. (2022) and derived the exact scale-by-scale budget equations for $\left\langle{\delta q^{2}}\right\rangle$, $\left\langle{\delta\omega^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, for polymeric turbulent flows. The equations are general and remain valid for a generic inhomogeneous and anisotropic turbulent flow as well. They capture the mechanisms of production, transfer and dissipation in the combined space of scales and positions. Compared to purely Newtonian flows, the polymer/fluid interaction introduces additional sink/source process and an alternative scale-space transfer for the three quantities. The newly derived equations have been then tailored to homogeneous and isotropic turbulence for the present problem.
@@
For $De\geq 1$ polymers effectively modify the cascades of energy and helicity, the two inviscid invariants of the $3D$ Navier–Stokes equations. Both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ are injected in the system at the largest scales and transferred to the small scales by two distinct transfer processes driven by fluid inertia and fluid/polymer interactions, where they are dissipated away. At large scales, the nonlinear cascade dominates, while at small scales the polymer-driven transfer takes over. The cross-over scale between the two dominant processes agrees fairly well $r_{p}^{*}$ at which $De(r_{p}^{*})\equiv\tau_{p}/\tau_{f}(r_{p}^{*})=1$, where $\tau_{p}$ and $\tau_{f}(r)$ are the polymer relaxation time and the characteristic turnover time of eddies at scale $r$. When $r<r_{p}^{*}$ and $\tau_{p}>\tau_{f}(r)$ $(De(r)>1)$ polymers effectively interact with the $r$-fluctuations of the carrier flow. The coexistence of the two transfer mechanisms leads to a multiscaling behaviour of the velocity and helicity structure functions. At small $De$ and large $r$, $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ exhibit the same power law predicted by the Kolmogorov theory, i.e. $\left\langle{\delta q^{2}}\right\rangle\sim\left\langle{\delta h}\right\rangle\sim r^{2/3}$ with $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle\approx 0.75$, consistently with helicity behaving as a passive scalar in HIT (Borue & Orszag, 1997). For $De\geq 1$ and at smaller scales, instead, both $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ deviate from the Kolmogorov predictions, with the latter showing a steeper slope. For $De=1$ we measure $\left\langle{\delta q^{2}}\right\rangle\sim r^{1.3}$ (in agreement with Zhang et al., 2021) and $\left\langle{\delta h}\right\rangle\sim r^{0.9}$. Accordingly, we observe that compared to the purely Newtonian case, the relative helicity $r\left\langle{\delta h}\right\rangle/2\left\langle{\delta q^{2}}\right\rangle$ increases with $De$ at all scales, indicating that polymers favour events that break mirror symmetry at small scales.
@@
A closer look at the energy fluxes reveals that polymers deplete the nonlinear cascade by weakening both the direct and inverse extreme events of both cascades, leading to a less skewed flux distribution as $De$ increases. For $r\gtrapprox r_{p}^{*}$, the nonlinear flux distribution collapses reasonably well for all $De$ once the quantities are normalised with their standard deviation. Moving to the polymeric flux, we observe that the amount of energy carried by the polymer-driven cascade increases with $De$ at all scales. Similar to the nonlinear flux, on average the polymeric flux transfers energy from larger to smaller scales, but local events with intense backscatter exist which transfer energy from small to large scales. However, the probability of very intense inverse transfer decreases as $De$ increases.
@@
Unlike energy and helicity, enstrophy is not an inviscid invariant of the $3D$ Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, which is related to the enstrophy produced by vortex stretching-like processes (Davidson, 2004). The conventional notion of cascade used for energy and helicity, thus, does not apply in this case, as most of the enstrophy is directly generated at the small scales. However, vortex stretching is active in the inertial range which results in an increasing rate of transfer of $\left\langle{\delta\omega^{2}}\right\rangle$ at small $r$, being maximum at the end of the inertial range. For $De\geq 1$, the transfer via the polymeric route is comparable with the fluid nonlinear contribution and, unlike for $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$, there is no transition with the two mechanisms coexisting in the entire inertial range. Overall, the net effect of polymers is to largely weaken vortex stretching, with fluid-polymer interactions becoming the primary enstrophy source at high $De$. The underlying local flow topology further reveals that vortex stretching modulation is a result of polymers promoting events with a two-dimensional state of straining, like shear and planar extensional flows. Accordingly, we observe that polymers favour events where rotation and stretching are equally strong (being typical of vortex sheets and shear layers), rather than those dominated by rigid rotation.
@@ Acknowledgments
The authors acknowledge the computer time provided by the Scientific Computing and Data Analysis section of Research Support Division at OIST, and by HPCI, under the Research Project grants hp210269, hp220099, hp230018, hp250021 and hp250035.
@@
Appendix A Energy spectra
Figure 21: Compensated energy spectra $k^{2.3}E(k)$ for $1\leq De\leq 9$
Figure 21 shows the scale-compensated energy spectra, $k^{2.3}E(k)$, for $1\leq De\leq 9$. In all cases, the compensated spectra display a well-defined plateau over an intermediate range of scales where polymer dynamics dominate, confirming the emergence of the $E(k)\sim k^{-2.3}$ scaling. The extent of this plateau is maximal at $De=1$ and gradually decreases with increasing $De$, consistent with the progressive decoupling between the fluid and polymeric phases.
It is worth stressing that this elastic self-similar behaviour is more clearly discernible in Fourier space ($E(k)$) rather than in physical space ($S_{2}(r)$), and the transition between regimes becomes evident earlier in the spectral domain. This can be attributed to the fact that second-order structure function is, loosely speaking, an integrated representation of the energy spectrum. As a result, the $S_{2}$ scaling exponent observed at a given scale $r$ reflects the cumulative contribution from all wavenumbers larger than $k\gtrsim 2\pi/r$. Rosti et al. (2023) showed that the range over which the spectrum follows the $k^{-2.3}$ scaling becomes narrower for $De>1$, while the classical $k^{-5/3}$ inertial range correspondingly expands. Based on this reasoning, the second-order structure function exhibits a combination of the $1.3$ and classical $2/3$ scalings at intermediate $De>1$. The classical classical $2/3$ scaling is fully recovered only at sufficiently large $De$, where the inertia-dirven cascade dominates across all intermediate scales.